OpenStudy (anonymous):

What is the simplified form of the quantity (15xy^2/x^2+5x+6)/(5x^2y/2x^2+7x+3) ?

OpenStudy (anonymous):

I believe it's the one I bubbled but I'm not sure

OpenStudy (kropot72):

The expression can be rewritten by inverting the denominator to give: \[\large \frac{15xy ^{2}}{x ^{2}+5x+6} \times \frac{2x ^{2}+7x+3}{5x ^{2}y}\] Do you understand this step?

OpenStudy (anonymous):

Yes I have that written down; I converted x^2 + 5x + 6 into (x + 2)(x + 3) And I also converted the 2x^2 + 7x + 3 into (2x + 1)(x + 3) \[\frac{ 15y }{ (x + 2)(x + 3)} \times \frac{ (2x + 1)(x + 3) }{ 5x }\]

OpenStudy (anonymous):

I cancelled out the exponents on 15xy^2 and 5x^2y

OpenStudy (anonymous):

Am I supposed to do that?

OpenStudy (kropot72):

You are correct so far. Now you can cancel (x + 3) in the denominator of the first term and in the numerator of the second term. Then divide 5x by 5 and divide 15y by 5.

OpenStudy (anonymous):

Is that because the GCF of both 5x and 15y is 5?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

\[\frac{ 3y }{ (x + 2) }\times \frac{ (2x + 1) }{ x }\] Like this?

OpenStudy (kropot72):

Good work! So basically you have finished. Which answer option have you got?

OpenStudy (anonymous):

\[\frac{ 3y(2x + 1) }{ x(x + 2) }\] Oh okay so I did get it right. Thanks!

OpenStudy (kropot72):

You're welcome :)